OpenStudy (anonymous):

• F is the force • m is the mass • t is the time in seconds a. Solve F= 27.2*M/ T for m. Please show your work. (6 points) b. Use your new formula to find the mass of the automobile in kg for each value of force and time given in the table below. Round your answer to the nearest whole number. (4 points) Force (Newtons) Mass (kg) Time (sec) 11000 ? 8 9520 ? 10 12000 ? 7 11500 ? 9

OpenStudy (anonymous):

For a. Solve for m; F= 27.2*M/ T Multiply by T on both sides T*F=27.2*M Divide by 27.2 on both sides T*F/27.2=M

OpenStudy (anonymous):

Force (Newtons) Mass (kg) Time (sec) 11000 ? 8 9520 ? 10 12000 ? 7 11500 ? 9

OpenStudy (anonymous):

find the mass

OpenStudy (anonymous):

Well use T*F/27.2=M i.e. 8*11000/27.2=M

OpenStudy (anonymous):

M=3235.3

OpenStudy (anonymous):

no that not it look at the chart

OpenStudy (anonymous):

M=3235kg for the first one.